Use De Moivre’s theorem to simplify the following (a) (cos π/5+i sin π/5)10, (b) (cos π/9+i sin π/9)−3, (c) {cos(−π/6) + i sin(−π/6)}−4,
1
Expert's answer
2020-05-25T20:29:17-0400

De Moivre’s Theorem:

"(\cos \theta +i \sin \theta )^n=\cos n\theta +i \sin n\theta, \ \" for all integers "n" .


a) "(\\cos\u03c0\/5+i\\sin\u2061\u03c0 \/5)^{10}=\\cos 2\\pi +i \\sin 2\\pi=1+i \\times 0=1"

b) "(\\cos\u2061\u03c0\/9+i\\sin\u2061\u03c0\/9)^{-3}=\n\\cos\n\u2061\n(\n\u2212\n\u03c0\n\/\n3\n)+i\n\\sin\n\u2061\n(\n\u2212\n\u03c0\n\/\n3\n)\n=\n1\/2+i (-\\sqrt 3 \/2)"

c) "(\cos(−π/6)+i\sin(−π/6)) 
^{−4}=\cos 2\pi /3+i \sin 2\pi/3=-1/2+i \sqrt 3 /2"